patty is 6 years old that marice. in 4 years, she will be twice as old as marice was 5 years ago. Find thier ages now
Answer by lwsshak3(11628)

patty is 6 years older that marice. in 4 years, she will be twice as old as marice was 5 years ago. Find their ages now
***
let x=marice's age
x+6=patty's age
..
In 4 years:
x+6+4=x+10=patty's age
..
5 years ago:
x-5=marice's age
..
(x+10)=2(x-5)
x+10=2x-10
x=20
x+6=26
marice's age=20
patty's age=26